Find a6 for an=8(1/2)^n, n=1,2,3,4,... I have no idea how to do this problem. Please help me.
Answer by fractalier(6550)

Well to find the sixth term of the sequence, a6, you merely plug the 6 in for n in the expression
an = 8 * (1/2)^n
a6 = 8 * (1/2)^6 = 8 * (1/64) = 1/8